An independent-measures study comparing two treatment conditions produces a t statistic with df = 18. If the two samples are the same size, how many participants were in each of the samples?

a) 9

b) 10

c) 19

d) 20
Answer #1

df = n1+n2-2

==> df = n+n-2 = 2n-2

==> n = (18+2)/2 = 10

10 samples in each

b) 10